# Gravitational wave probes of particle dark matter: a review

### v3

@@ II.3.2 Changes to neutron-star moment of inertia
Fluctuations in fundamental constants could also cause other effects, such as changing the moment of inertia of neutron stars Kaplan et al. (2022), or changing the times measured cesium clocks Kaplan et al. (2022). These effects will be relevant when we describe how ultralight dark matter could be detected in pulsar timing arrays. To study the former, we must also consider interactions in QCD Damour and Donoghue (2010); Kaplan et al. (2022):
@@
where $\beta_{3}$ is the QCD beta function, $g_{3}$ is the QCD gauge coupling, $G_{\mu\nu}^{A}$ is the gluon field strength tensor, $\gamma_{q}$ are the light quark anomalous dimensions, $\Lambda=M_{\text{Pl}}/\sqrt{4\pi}$, and $d_{g}$, and $d_{q}$ are the dilaton coupling constants to the gluon and the quarks, respectively. Reading off the Lagrangian, and noting the difference from the standard model, the change in masses of quarks is:
$\frac{\delta m_{q}}{m_{q}}$ $=\frac{d_{q}}{\Lambda}\phi\,,$ (32)
while in nucleon masses:
@@
where $C_{n}=0.048$ Damour and Donoghue (2010), $m_{p}$ is proton mass, $m_{n}$ is the neutron mass, and the symmetric combination of the quark mass couplings has been defined to be:
$d_{\hat{m}}\equiv\frac{d_{u}m_{u}+d_{d}m_{d}}{m_{u}+m_{d}}\,.$ (34)
Moving back to the neutron star, conservation of angular momentum requires that the fluctuations in the spin frequency $\delta\omega$ relate to changes in the pulsar’s moment of inertia $\delta I$ as:
@@
where $I_{0}$ is the unperturbed moment of inertia and $\omega_{0}$ is the unperturbed angular frequency of the neutron star. Based on the above discussion, DM interactions can cause changes in the masses of fundamental particles, which, in turn, cause time-dependent oscillations in the moment of inertia. In the case of neutron stars, the mass of the neutron will oscillate, and thus, indirectly, the mass of the neutron star itself will change as well.
Assuming a simplistic model of a non-rotating, spherically symmetric neutron star composed only of neutrons with a polytropic equation-of-state, the radius and moment of inertia can be written as Shapiro and Teukolsky (1983):
@@
where $R$ and $M_{0}$ are the radius and unperturbed mass of the neutron star. Solving explicitly the Tolman–Oppenheimer–Volkoff (TOV) equations for a realistic equation-of-state would cause $\mathcal{O}(1)$ deviations from the simple scalings. These “unknown” scalings can be parameterized in terms of $\eta$ and $\delta\eta$
@@
which are set to $\eta=1/3$ and $\delta\eta=-16/3$ in line with the simplistic model, and to illustrate the concept.
Because of the changes in the pulsar’s moment of inertia, the pulsar’s spin frequency will change. For pulsar timing array experiments, the measured quantities are the times of arrivals of pulses from each pulsar, which will change in response to the oscillating moment of inertia. In practice, the observable here is the residual $h$ of the times of arrivals of pulses from the neutron star, which is given by:
[…]
@@
$\frac{\delta M}{M_{0}}$ $=\sum_{f\in\{e,\mu,p,n\}}Y_{f}\frac{m_{f}}{m_{n}}\frac{\delta m_{f}}{m_{f}}\,,$ (40)
where $\delta M$ is the change in the pulsar’s mass. The sum is over the different types of particles present in the pulsar: electrons, muons, protons and neutrons, $Y_{f}\equiv N_{f}/(N_{n}+N_{p})$ is a weighting function that gives the number of $f$ particles relative to the sum of protons and neutrons (which dominate the pulsar’s mass). For neutron star observed by pulsar timing arrays, the following is assumed for $Y_{f}$: $Y_{n}=0.9$, $Y_{p}=Y_{e}=0.1$ and $Y_{\mu}=0.05$ Bell et al. (2019); Kaplan et al. (2022).
@@
$h(t)=$ $\frac{\sqrt{2\rho_{\phi}}}{m_{\rm DM}^{2}\Lambda}\left(\vec{y}\cdot\vec{d}\,\right)\hat{\phi}_{P}\sin{\big{(}m_{\rm DM}t+\gamma(\vec{x}_{P})\big{)}}\,,$ (41)
@@
[…]
This type of DM interaction would leave imprints on pulsar timing array experiments Porayko et al. (2018); Smarra et al. (2023); Afzal et al. (2023), which will be discussed later.
@@ II.3.3 Changes to atomic clock times
Changes in fundamental constants would also affect the way in which atomic clocks measure time. Since pulsar timing array experiments reference their measurements of the times of arrivals of pulses to terrestrial time (TT) using atomic clocks composed primarily of cesium, any oscillation of the fundamental constants will affect the atomic clocks, and thus the times measured Flambaum et al. (2004). Likewise, a shift in the frequency of the atomic clock will give rise to an apparent shift in the measured pulsar spin frequency, given by Eq. 39.
@@
where $Z$ is the nuclear charge, $F_{\rm rel}(Z\alpha)$ is a relativistic correction to an atom’s energy levels, $\mu$ is the nuclear magnetic moment, and $\zeta=1$ if the clocks that have hyperfine transitions (the case of cesium), and $\zeta=0$ for optical transitions [^2] This equation arises from solving the non-relativistic Schrodinger equation for the valence electron wavefunction and evaluating it at the nucleus of the atom.. The Casmir effect $F_{\rm rel}(Z\alpha)$ is a correction that comes from solving the relativistic wave equation for an electron near the nucleus
The fluctuations in fundamental constants will cause fluctuations in the frequencies of atomic clocks composed of atoms $A$:
$\begin{split}\frac{\delta f_{A}}{f_{A}}\simeq&\Bigg{[}\frac{\delta m_{e}}{m_{e}}+(4+K_{A})\frac{\delta\alpha}{\alpha}\\ &+\zeta\Bigg{(}\frac{\delta m_{e}}{m_{e}}+C_{A}\sum_{q=u,d}\frac{\delta m_{q}}{m_{q}}-\frac{\delta m_{p}}{m_{p}}\Bigg{)}\Bigg{]}\,,\end{split}$ (44)
where the derivative of the Casmir effect term is $\delta F_{\rm rel}/F_{\rm rel}=K_{A}$, and the derivative of the nuclear magnetic moment is $\delta\mu/\mu=C_{A}\,\delta m_{q}/m_{q}$. For Cesium, $K_{A}=0.83$ depends nontrivially on $Z$ and $\alpha$, and $C_{A}=0.110$ Flambaum et al. (2004). $m_{q}$ is a quark mass, and the index $q$ runs over up and down quarks.
Finally, by substituting the aforementioned equations for fluctuations of the fundamental constants into Eq. 44, and then further plugging this result into Eq. 39, the timing residuals induced by the ultralight dark-matter are :
$h(t)=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}^{2}\Lambda}\left(\vec{y}\cdot\vec{d}\right)\hat{\phi}_{E}\sin{\big{(}m_{\rm DM}t+\gamma\big{)}}\,.$ (45)
@@
where $\xi_{A}\equiv 4+K_{A}$, and $\hat{C}_{A}=C_{A}(m_{u}+m_{d})^{2}/2m_{u}m_{d}$. This type of DM interaction would leave imprints on pulsar timing array experiments Porayko et al. (2018); Smarra et al. (2023); Afzal et al. (2023), which will be discussed later.
@@ II.4 Vector dark matter
DM could be spin-1, i.e. vectors, which we refer to as “dark photons”. Dark photons could explain the entirety of the relic abundance of DM, which could arise from the misalignment mechanism Nelson and Scholtz (2011); Arias et al. (2012); Graham et al. (2016), parametric resonance or the tachyonic instability of a scalar field Agrawal et al. (2020); Co et al. (2019); Bastero-Gil et al. (2019); Dror et al. (2019), or from cosmic string network decays Long and Wang (2019). They could couple to standard model particles, either to baryon or baryon-lepton number. In the interferometers, these interactions would occur everywhere, but would be detectable in the four interferometer mirrors. The dark photons [^3] These are different dark photons than those that kinetically mix with the ordinary photon, which will be discussed later in the context of luminous superradiance Siemonsen et al. (2023a). would cause a “dark” force on the mirrors, causing quasi-sinusoidal oscillations Pierce et al. (2018); Miller et al. (2021).
Dark photons can be described analogously to ordinary photons, i.e. with a four-vector potential, which can be written as:
[…]
where $(A_{0})_{\mu}$ is the four-amplitude of $A_{\mu}$ $\Upsilon$ is a random phase, and $\vec{x}$ is the position at which $A_{\mu}$ is measured. The index $\mu={0,1,2,3}$ refers to both time and spatial components.
@@
$\frac{(A_{0})_{0}}{|\vec{A}_{0}|}=\frac{v_{0}}{c}\simeq 7.667\times 10^{-4},$ (48)
@@
where $|\vec{A}_{0}|=\left(\frac{\hbar\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}c^{2}}\frac{1}{\sqrt{\epsilon_{0}}}\right)$ is the magnitude of the spatial components of $A_{\mu}$, normalized by the present DM energy density of the universe, and $\epsilon_{0}$ is the permittivity of vacuum. In Eq. 48, we see that the dark scalar potential is suppressed by about three orders of magnitude compared to the the dark three-vector potential. Therefore, the time-component of the four-vector potential can be safely neglected, leaving only the three-vector potential:
$\vec{A}=\vec{A}_{0}\sin\left(\omega t-\vec{k}\cdot\vec{x}+\Upsilon\right),$ (49)
@@
The Lagrangian $\mathcal{L}$ that characterizes the dark photon coupling to a number current density $J^{\mu}$ of baryons or baryons minus leptons is:
$\mathcal{L}=-\frac{1}{4\mu_{0}}F^{\mu\nu}F_{\mu\nu}+\frac{1}{2\mu_{0}}\left(\frac{m_{\rm DM}c}{\hbar}\right)^{2}A^{\mu}A_{\mu}-\epsilon eJ^{\mu}A_{\mu},\$ (50)
@@
where $\mu_{0}$ is the magnetic permeability in vacuum and $\epsilon$ is the strength of the particle/dark photon coupling normalized by the electromagnetic coupling constant.
Using Eq. 47, the “dark” electric and magnetic fields can be derived:
$\vec{E}$ $=\partial_{0}\vec{A}-\vec{\nabla}A_{0}\simeq\omega\vec{A}_{0}\cos(\omega t-\vec{k}\cdot\vec{x}+\phi),$ (51)
$\vec{B}$ =×$=\vec{\nabla}\times\vec{A}=-\vec{k}\times\vec{A}_{0}\cos(\omega t-\vec{k}\cdot\vec{x}+\phi),$ (52)
@@
and we note that the electric field is much stronger than the magnetic one:
$\frac{|\vec{E}|}{|c\vec{B}|}\sim\frac{\omega}{|\vec{k}|}=\frac{c^{2}}{|\vec{v}|}\sim 10^{3}c,$ (53)
@@
which implies that we need to only consider the electric field. This dark electric field causes the test masses to oscillate at a quasi-fixed frequency given by the dark photon mass. The acceleration of a given mirror can thus be derived Guo et al. (2019); Pierce et al. (2018):
$\vec{a}_{j}(t,\vec{x}_{j})$ $=\frac{\vec{F}_{j}(t,\vec{x}_{j})}{M_{j}}\simeq\epsilon e\frac{Q_{D,j}}{M_{j}}\omega|\vec{A}_{0}|\hat{A}\cos(\omega t-\vec{k}\cdot\vec{x}_{j}+\phi),$
$\epsilon^{2}$ $=\frac{\alpha_{\text{DP}}}{\alpha}$ (54)
@@
where $\alpha_{\text{DP}}$ is the dark photon fine structure constant, $\epsilon^{2}$ is the ratio between the dark photon fine structure constant and the electromagnetic fine structure constant $\alpha$, and $Q_{D,j}$ is the total charge in the $j$th mirror of mass $M_{j}$. $\hat{A}$ is a unit vector, and $\phi$ is a function of the spatial coordinates. If dark photons couple to the baryon number, $q_{j}$ is the number of protons and neutrons in each mirror; if they couple to the difference between the baryon and lepton numbers, $q_{j}$ is the number of neutrons in each mirror. Each mirror is positioned differently with respect to the incoming dark photon DM field, and thus is accelerated at a slightly different amount over time. Thus, the interferometer experiences a measurable differential acceleration, and thus a differential strain.
@@
Assuming that all mirrors are the same, integrating Eq. 54 twice over time, and averaging over random polarization and propagation directions, the strain on the interferometers caused by a dark photon DM signal is Pierce et al. (2018):
$\sqrt{\langle h^{2}_{D}\rangle}$ $=C\frac{Q_{D}}{M}\frac{\hbar e}{c^{4}\sqrt{\epsilon_{0}}}\sqrt{2\rho_{\text{DM}}}v_{0}\frac{\epsilon}{f_{0}},$
@@
where $C=\sqrt{2}/3$ (in LIGO) is a geometrical factor obtained by averaging over all possible dark photon propagation and polarization directions, and different detector geometricians (the calculation for $C$ is shown in the appendix of Pierce et al. (2018) and in Morisaki et al. (2021)).
A second strain results due to the so-called “finite light travel time effect, in which the mirrors in the interferometers have moved in the time that it takes the light to reach them from the beamsplitter. This strain can actually be larger than that given in Eq. 55:
@@
The total strain is: $\langle h_{{\rm total}}^{2}\rangle=\langle h_{D}^{2}\rangle+\langle h_{C}^{2}\rangle$.

### v4

@@ II.3.2 Changes to neutron-star moment of inertia
Beyond laboratory-based detectors, dilaton-induced fluctuations in fundamental constants could also have observable consequences in astrophysical systems. For example, they could alter moment of inertia of neutron stars Kaplan et al. (2022), which would produce deviations in the arrival time of pulses, to be discussed in Section III.3. In order to quantify the extent to which the neutron star moment of inertia will change, we must know how protons and neutrons will change in response to dilatons. This motivates us to consider the following interactions in QCD of quarks in the presence of dilatons Damour and Donoghue (2010); Kaplan et al. (2022):
@@
where $\beta_{3}$ is the QCD beta function, $g_{3}$ is the QCD gauge coupling, $G_{\mu\nu}^{A}$ is the gluon field strength tensor, $\gamma_{q}$ are the light quark anomalous dimensions, $\Lambda=M_{\text{Pl}}/\sqrt{4\pi}$, $m_{q}$ is the quark mass, and $d_{g}$ and $d_{q}$ are the dilaton coupling constants to the gluon and the quarks, respectively.
Similarly to Section II.3, we can write the standard-model Lagrangian for quarks:
$\mathcal{L}_{\text{quark}}^{\text{SM}}=-\sum_{q}m_{q}\bar{q}q,$ (30)
which, upon comparing with Eq. 29, allows us to read off the change in quark mass as
$\frac{\delta m_{q}}{m_{q}}$ $=\frac{d_{q}}{\Lambda}\phi.$ (31)
The shift in nucleon masses can then be obtained by noting that nucleons are composite objects whose masses arise mostly from QCD binding energy, with a smaller contribution from the quark masses. The dilaton couples to both the quark masses and the gluon field strength, so the fractional change in a nucleon mass is a weighted sum of these effects Damour and Donoghue (2010); Kaplan et al. (2022):
@@
where $C_{n}$ accounts for the fraction of the nucleon mass due to the light quarks, and
$d_{\hat{m}}=\frac{m_{u}d_{u}+m_{d}d_{d}}{m_{u}+m_{d}}$ (33)
is the effective quark coupling. In other words, the nucleon mass responds both to the direct change in quark masses induced by the dilaton and to the change in the gluon’s binding energy, allowing us to propagate the fundamental dilaton couplings to macroscopic observables such as the neutron-star moment of inertia.
Eqs. 31 and 32 show how dilaton DM will alter the masses of quarks and nucleons, which are the primary constitutions of neutron stars. Thus, these changed masses will impact the rotational frequency and moment of inertia of neutron stars, which will affect the times of arrivals (ToAs) of pulses that come from them. To calculate these effects, let us recall that conservation of angular momentum requires that the fluctuations in the spin frequency $\delta\omega$ relate to changes in the pulsar’s moment of inertia $\delta I$ as:
@@
where $I_{0}$ is the unperturbed moment of inertia and $\omega_{0}$ is the unperturbed angular frequency of the neutron star. Assuming a simplistic model of a non-rotating, spherically symmetric neutron star composed only of neutrons with a polytropic equation-of-state, the radius and moment of inertia can be written as Shapiro and Teukolsky (1983):
@@
where $R$ and $M_{0}$ are the radius and unperturbed mass of the neutron star, respectively. Plugging in Eqs. 35 and 36 into Eq. 34, we can derive how the moment of inertia changes as a function of the neutron-star mass and neutron mass:
@@
where $\delta M$ is the change in the pulsar’s mass, and $\eta=1/3$ and $\delta\eta=-16/3$ in this simplistic model [^2] Solving explicitly the Tolman–Oppenheimer–Volkoff (TOV) equations for a realistic equation-of-state would cause $\mathcal{O}(1)$ deviations from the simple scalings. These “unknown” scalings can be parameterized in terms of $\eta$ and $\delta\eta$.
Because of the changes in the pulsar’s moment of inertia, the pulsar’s spin frequency will change. For pulsar timing array experiments, the measured quantities are the ToAs of pulses from each pulsar, which will change in response to the oscillating moment of inertia. In practice, the observable here is the residual $h$ of the ToAs of pulses from the neutron star, which is given by:
$h=-\int\frac{\delta\omega}{\omega_{0}}dt=\int\frac{\delta I}{I_{0}}dt\,.$ (38)
@@
$\frac{\delta M}{M_{0}}$ $=\sum_{f\in\{e,\mu,p,n\}}Y_{f}\frac{m_{f}}{m_{n}}\frac{\delta m_{f}}{m_{f}}\,.$ (39)
The sum is over the different types of particles present in the pulsar: electrons, muons, protons and neutrons, $Y_{f}\equiv N_{f}/(N_{n}+N_{p})$ is a weighting function that gives the number of $f$ particles relative to the sum of protons $N_{p}$ and neutrons $N_{n}$ (which dominate the pulsar’s mass). For neutron stars observed by pulsar timing arrays, the following is assumed for $Y_{f}$: $Y_{n}=0.9$, $Y_{p}=Y_{e}=0.1$ and $Y_{\mu}=0.05$ Bell et al. (2019); Kaplan et al. (2022).
@@
$h(t)=$ $\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}^{2}\Lambda}\left(\vec{y}\cdot\vec{d}\,\right)\hat{\phi}_{P}\sin{\big{(}m_{\rm DM}t+\gamma(\vec{x}_{P})\big{)}}\,$, (40)
@@
$\begin{split}\{y_{g},\,y_{\hat{m}},\,y_{\mu},\,y_{e}\}=&\,\eta\left\{1,\,C_{n},\,6\times 10^{-3},\,5\times 10^{-5}\right\}\\ +&\delta\eta\left\{1,\,C_{n},\,0,\,0\right\}\,.\end{split}$ (41)
Using the above formalism, dilatons have been constrained with recent pulsar timing array experiments Porayko et al. (2018); Smarra et al. (2023); Afzal et al. (2023), which will be extensively discussed in Section III.3.
@@ II.3.3 Changes to atomic clock times
Changes in fundamental constants would also affect the way in which atomic clocks measure time. Since pulsar timing array experiments reference their measurements of the ToAs of pulses to terrestrial time (TT) using atomic clocks composed primarily of cesium, any oscillation of the fundamental constants will affect the atomic clocks, and thus the times measured Flambaum et al. (2004). Likewise, a shift in the frequency of the atomic clock will give rise to an apparent shift in the measured pulsar spin frequency, given by Eq. 38.
@@
where $Z$ is the nuclear charge, $F_{\rm rel}(Z\alpha)$ is a relativistic correction to an atom’s energy levels, $\mu$ is the nuclear magnetic moment, and $\zeta=1$ if the clocks that have hyperfine transitions (the case of cesium), and $\zeta=0$ for optical transitions [^3] This equation arises from solving the non-relativistic Schrodinger equation for the valence electron wavefunction and evaluating it at the nucleus of the atom..
The fluctuations in fundamental constants will cause modulations in the frequencies of atomic clocks composed of atoms $A$. Using previously derived expressions for variations in the electron mass, fine-structure constant, and quark and nucleon masses, the fractional frequency shift can be obtained Flambaum et al. (2004); Kaplan et al. (2022):
[…]
where $\delta F_{\rm rel}/F_{\rm rel}=K_{A}$, and the derivative of the nuclear magnetic moment is $\delta\mu/\mu=C_{A}\,\delta m_{q}/m_{q}$. For Cesium, $K_{A}=0.83$ depends on $Z$ and $\alpha$, and $C_{A}=0.110$ Flambaum et al. (2004). The index $q$ runs over up and down quarks.
Finally, by substituting the aforementioned equations for fluctuations of the fundamental constants (Eqs. 21, 22, 31 and 32) into Eq. 43, and then further plugging this result into Eq. 38, the timing residuals induced by the ultralight dark-matter can be obtained:
$h(t)=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}^{2}\Lambda}\left(\vec{y}\cdot\vec{d}\right)\hat{\phi}_{E}\sin{\big{(}m_{\rm DM}t+\gamma(\vec{x}_{E})\big{)}}\,.$ (44)
@@
where $\xi_{A}\equiv 4+K_{A}$, and $\hat{C}_{A}=C_{A}(m_{u}+m_{d})^{2}/2m_{u}m_{d}$. This type of DM interaction would leave imprints on pulsar timing array experiments Porayko et al. (2018); Smarra et al. (2023); Afzal et al. (2023), which will be discussed in Section III.3.
We would like to draw a parallel between Eqs. 40 and 45. Both equations have a similar form, though the underlying physics is different, as indicated by the phase factors: the former depends on the phase at the neutron star, while the latter depends on the phase on earth (at the atomic clocks). Of course, the sensitivity parameters of the search are different, which encode which couplings that each effect depends on. These two effects highlight an important point that will be made throughout this review: while the physical signatures of different kinds of DM are different, often the impact on the signal (the residuals in pulsar timing arrays, the phase shifts in ground- and spaced-based interferometers) are similar, and only the interpretation of the lack of a detection is different.
@@ II.4 Vector dark matter
So far, we have considered explain how spin-0 particles would affect GW interferometers; however, DM can also have nonzero spins. Here, we consider a model for DM that has spin-1, i.e. vectors, which we refer to as “dark photons”. Like dilatons, dark photons could explain the entirety of the relic abundance of DM, which could arise from the misalignment mechanism Nelson and Scholtz (2011); Arias et al. (2012); Graham et al. (2016), parametric resonance or the tachyonic instability of a scalar field Agrawal et al. (2020); Co et al. (2019); Bastero-Gil et al. (2019); Dror et al. (2019), or from cosmic string network decays Long and Wang (2019). They could couple to standard-model particles, either to baryon or baryon-lepton number. In the interferometers, these interactions would occur everywhere, but would be most pronounced in the four interferometer test masses in the main resonant cavities of the instrument. The dark photons [^4] These are different dark photons than those that kinetically mix with the ordinary photon, which will be discussed later in the context of luminous superradiance Siemonsen et al. (2023a). would cause a “dark” force on the mirrors, causing quasi-sinusoidal oscillations Pierce et al. (2018); Miller et al. (2021a), as will be shown below.
To derive the impact of dark photons on GW interferometers, we will begin by describing dark photons analogously to ordinary photons, i.e. with a four-vector potential:
$A_{\mu}(t,\vec{x})=(A_{0})_{\mu}\sin(m_{\rm DM}t-\vec{k}\cdot\vec{x}+\Upsilon),$ (46)
where $(A_{0})_{\mu}$ is the four-amplitude of $A_{\mu}$ $\Upsilon$ is a random phase, $\vec{k}$ is the wavevector, and $\vec{x}$ is the position at which $A_{\mu}$ is measured. The index $\mu={0,1,2,3}$ refers to both time and spatial components.
@@
$\frac{(A_{0})_{0}}{|\vec{A}_{0}|}=v_{0}\simeq 7.667\times 10^{-4},$ (47)
@@
where $|\vec{A}_{0}|=\left(\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}}\right)$ is the magnitude of the spatial components of $A_{\mu}$, normalized by the present DM energy density of the universe. In Eq. 47, we see that the dark scalar potential is suppressed by about three orders of magnitude compared to the the dark three-vector potential. Therefore, the time-component of the four-vector potential can be safely neglected, leaving only the three-vector potential:
$\vec{A}=\vec{A}_{0}\sin\left(m_{\rm DM}t-\vec{k}\cdot\vec{x}+\Upsilon\right),$ (48)
@@
With the four-potential, we can write the Lagrangian $\mathcal{L}$ that characterizes the dark photon coupling to a number current density $J^{\mu}$ of baryons or baryons minus leptons:
$\mathcal{L}=-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\frac{1}{2}m_{\rm DM}^{2}A^{\mu}A_{\mu}-\epsilon eJ^{\mu}A_{\mu},\\$ (49)
@@
where $\epsilon$ is the strength of the particle/dark photon coupling normalized by the electromagnetic coupling constant.
Using Eqs. 46 and 48, we can compute the “dark” electric and magnetic fields, analogously to electromagnetism, which will subsequently allow us to compute the “dark” force:
$\vec{E}$ $=\partial_{0}\vec{A}-\vec{\nabla}A_{0}\simeq m_{\rm DM}\vec{A}_{0}\cos(m_{\rm DM}t-\vec{k}\cdot\vec{x}+\Upsilon),$ (50)
$\vec{B}$ $=\vec{\nabla}\times\vec{A}=-\vec{k}\times\vec{A}_{0}\cos(m_{\rm DM}t-\vec{k}\cdot\vec{x}+\Upsilon),$ (51)
@@
It is important to understand the relative magnitudes of the electric and magnetic fields. Comparing them, we see that the electric field is much stronger than the magnetic one:
$\frac{|\vec{E}|}{|\vec{B}|}\sim\frac{m_{\rm DM}}{|\vec{k}|}=\frac{1}{|\vec{v}|}\sim 10^{3},$ (52)
@@
which implies that we need to only consider the electric field. This dark electric field causes the test masses to oscillate at a quasi-fixed frequency given by the dark photon mass. The acceleration of a given mirror can thus be derived using Eqs. 49 and 48 Guo et al. (2019); Pierce et al. (2018):
$\vec{a}_{j}(t,\vec{x}_{j})$ $=\simeq\epsilon e\frac{Q_{D,j}}{M_{j}}m_{\rm DM}|\vec{A}_{0}|\hat{A}\cos(m_{\rm DM}t-\vec{k}\cdot\vec{x}_{j}+\Upsilon),$
@@
where $Q_{D,j}$ is the total charge in the $j$th mirror of mass $M_{j}$, and $\hat{A}$ is a unit vector of the vector potential. If dark photons couple to the baryon number, $q_{j}$ is the number of protons and neutrons in each mirror; if they couple to the difference between the baryon and lepton numbers, $q_{j}$ is the number of neutrons in each mirror. Each mirror is positioned differently with respect to the incoming dark photon DM field, and thus is accelerated at a slightly different amount over time. Thus, the interferometer experiences a measurable differential acceleration, and thus a differential strain.
@@
Assuming that all mirrors have the same charge-to-mass ratio, integrating LABEL:eqn:accel twice over time, and averaging over random polarization and propagation directions, the strain on the interferometers caused by a dark photon DM signal can be computed Pierce et al. (2018):
$\sqrt{\langle h^{2}_{D}\rangle}$ $=C\frac{Q}{M}e\sqrt{2\rho_{\text{DM}}}v_{0}\frac{\epsilon}{f_{0}},$
@@
where $C=\sqrt{2}/3$ (for an interferometer with two perpendicular arms) is a geometrical factor obtained by averaging over all possible dark photon propagation and polarization directions, and different detector geometries (the calculation for $C$ is shown in the appendix of Pierce et al. (2018) and in Morisaki et al. (2021)).
A second strain results due to the so-called “finite light travel time effect, in which the mirrors in the interferometers have moved in the time that it takes the light to reach them from the beamsplitter. This strain can actually be larger than that given in Eq. 54 and can be derived Morisaki et al. (2021):
@@
The total power is: $\langle h_{{\rm total}}^{2}\rangle=\langle h_{D}^{2}\rangle+\langle h_{C}^{2}\rangle$.
Both of these effects can be constrained by analyzing data from LIGO, Virgo and KAGRA and pulsar timing arrays, which will be discussed in Sections III.2 and III.3, though in different mass regimes.